OpenStudy (anonymous):

A statement Sn about the positive integers is given. Write statements S1, S2, and S3, and show that each of these statements is true. Show your work. Sn: 2 is a factor of n2 + 7n

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

@mathmath333 @Nnesha

OpenStudy (cwrw238):

Sn = n^2 + 7n so what is the value of S1?

OpenStudy (cwrw238):

just substitute 1 for n in the above formula and work it out

OpenStudy (mathmath333):

are u required to use mathmatical induction here

OpenStudy (mathmath333):

like in the previous problem

OpenStudy (anonymous):

the qustion dont ask for it @mathmath333

OpenStudy (mathmath333):

ok so sbustitue the values of \(n=1,2~and~3\) in \(n2 + 7n\)

OpenStudy (anonymous):

i am unable to post wait a sec i already did that

OpenStudy (anonymous):

S(n) : 2 is a factor of n^2 + 7n

OpenStudy (anonymous):

S(1) : 2 is a factor of 1^2 + 7*1 , S(2) : 2 is a factor of 2^2 + 7*2

OpenStudy (anonymous):

are S(1) and S(2) true

OpenStudy (anonymous):

i did this but my teacher said : Substitute values of 1, 2 and 3 for n and then work out the formula. Are each of these a multiple of two or even?

OpenStudy (mathmath333):

what is \(1^2 + 7*1=\)

OpenStudy (anonymous):

i guess i already did that?

OpenStudy (anonymous):

8

OpenStudy (mathmath333):

is \(2\) a factor of \(8\)

OpenStudy (anonymous):

then 53 and S(3) is 30

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

2*4=8

OpenStudy (anonymous):

8 and 30 is multiple of 30

OpenStudy (mathmath333):

so is it true for \(S_1\)

OpenStudy (anonymous):

S(2) : 2 is a factor of 2^2 + 7*2 =53 S(3): 2 is a factor of 3^2+7*3=30 are S(1) and S(3) is true

OpenStudy (anonymous):

am i correct? @mathmath333

OpenStudy (mathmath333):

what is \((2^2+7\times 2)\)

OpenStudy (anonymous):

oops i thought its square and cube

OpenStudy (anonymous):

18

OpenStudy (mathmath333):

is \(2\) a factor of \(18\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

2*4=8 2*9=18

OpenStudy (anonymous):

and 2*15=30?

OpenStudy (anonymous):

please reply soon i have to go teach my little bro he have revision exam tomorrow

OpenStudy (mathmath333):

its right

OpenStudy (anonymous):

thanks